OpenStudy (anonymous):

subtract -1.5-(-3.6)

OpenStudy (abb0t):

Two negatives = + Therefore, you have: -1.5 + 3.6 = 2.1 Think of it as money, you have $3.60 (+3.6) , but you owe someone $1.50 (-1.5)

OpenStudy (mathstudent55):

-1.5 - (-3.6) = -1.5 + 3.6 Minus negative 3.6 equals positive 3.6